In your chemistry lab you have a bottle of 1% acid and a bottle of 5% of acid. You need 100milliliters of 3%of acid for something. How many milliliters of each acid do you need to mix together
Answer by mananth(16949)

percent ---------------- quantity
1 ---------------- x
5 ---------------- 100-x
3.00% ---------------- 100
...
x+5(100-x)=3*100

x+500-5x =300
x-5x=300 -500
-4x-=-200
/-4
x=50 ml 1%
50ml 3%